# Full asymptotic expansions of the Humbert function $\Phi_{1}$

### v1

@@ Abstract
We derive full asymptotic expansions for the Humbert function $\Phi_{1}$ in different limiting regimes of its variables. Our derivation employs various asymptotic methods and relies on key transformation formulae established by Erdélyi (1940), and Tuan and Kalla (1987). The efficiency of our asymptotic results are also illustrated through two applications: (1) analytic continuations of Saran’s function $F_{M}$, and (2) two limits arising in the study of the $1D$ Glauber-Ising model. Finally, some promising directions for future research are highlighted.
@@ 1.  Introduction
Applying the Euler transformation [18, Eq. (15.8.1)]
@@
Moreover, the generalized hypergeometric function ${}_{p}F_{q}$ is defined by
@@ 2.1.  Regime (i): $x\to\infty$
Using the connection formula for ${}_{2}F_{1}$ ([18, Eq. (15.8.2)])
@@ 2.3.  Regime (iii): $y\to\infty$
Recall the estimate [11, Eq. (3.4)]: as $t\to\pm\infty$,
@@
We close the path of integration by the semi-circle $\mathfrak{C}$, that is parameterized by $s=(N+1/2)\mathrm{e}^{\mathrm{i}\theta}$ $(|\theta|\leqslant\pi/2)$, and then we let $N\to\infty$ through integer values. Recall the estimate [11, p. 7]
@@ 2.4.  Regime (iv): $x\to\infty,\,y\to\infty$
and from asymptotics of the Humbert function $\Psi_{1}$ for two large arguments (see [13, Theorems 1.2 and 1.3]). We formulate the precise asymptotic expansions in the following two theorems.
@@
$\Phi_{1}={}$ $\frac{\Gamma(c)\Gamma(b-a)}{\Gamma(b)\Gamma(c-a)}A_{1}^{*}(x,y)+\frac{\Gamma(c)\Gamma(a-b)}{\Gamma(a)\Gamma(c-a)}\Bigl{(}A_{2}^{*}(x,y)+A_{3}^{*}(x,y)\Bigr{)}$
$+\frac{\Gamma(c)}{\Gamma(a)}A_{4}^{*}(x,y)+\mathcal{O}\left(\left|y\right|^{-\Re(b)-w}\right)+\mathcal{O}\left(\left|y\right|^{\Re(a-b-c)-N}\mathrm{e}^{\Re(y)}\right),$
@@
$A_{2}^{*}(x,y)={}$ $\left(-\beta\right)^{b-a}\mathrm{e}^{-\beta}\left(1-x\right)^{-a}\sum_{k=1}^{M}\frac{\left(a\right)_{k}}{k!}\left(1-x\right)^{-k}\sum_{j=0}^{k-1}\frac{\left(c-a\right)_{j}\left(a-b\right)_{k-j}\left(a-c+1\right)_{k-j}}{j!\left(a\right)_{k-j}}\left(-\beta\right)^{j-k},$
$A_{3}^{*}(x,y)={}$ $\left(-\beta\right)^{b-a}\mathrm{e}^{-\beta}\left(1-x\right)^{-a}\sum_{k=0}^{M}\frac{\left(a\right)_{k}\left(c-a\right)_{k}}{\left.k!\right.^{2}}\,_{3}F_{3}\left[\begin{matrix}1,1-a,c-a+k\\ c-a,b-a+1,1+k\end{matrix};\beta\right]\left(1-x\right)^{-k},$
$A_{4}^{*}(x,y)={}$ $\beta^{a-c}\left(1-x\right)^{a-b-c}\mathrm{e}^{y}\sum_{k=0}^{N-1}a_{k}^{*}\left(x,y\right)y^{-k},$
@@
2.5.  Regime (vi): $x$ or $y$ small, $xy$ fixed
@@
It suffices to establish Assertion (1), as the proof of (2) follows analogously through application of the estimate in [13, Theorem 2.6(ii)]. Now assume that the conditions stated in (1) are valid.
@@
Estimate on $S_{2}$. For $n\geqslant N$, use the estimate [13, Theorem 2.5] to get
@@ 3.1.  Analytic continuations of $F_{M}$
Saran’s function $F_{M}$ is defined by ([19, Eq. (2.5)])
@@
which, in view of (1.3) and the estimate [11, Eq. (2.2)]
@@
which corrects Saran’s original form [20, p. 134, Eq. (4)]. Our results in Section 2.3 regarding asymptotics of $\Phi_{1}[y,zt]$ for large $t$ then provide a sufficient convergence condition for (3.3):
@@
The final continuation of $F_{M}$ follows from the following Mellin-Barnes integral, which is established by repeating the analysis in the proof of Theorem 2.3 and applying the asymptotic results of the Gauss hypergeometric function ${}_{2}F_{1}$ for large parameters in [11, Section 2.1].
@@ 3.2.  $1D$ Glauber-Ising model
which, by using the reduction formula [18, Eq. (15.4.3)]
@@ 4.  Discussion
which hold uniformly with respect to $n\in\mathbb{Z}_{\geqslant 0}$. More generally, it is necessary to extend the asymptotic expansions for ${}_{1}F_{1}[a;b;z]$ established in [24, Theorem 3.1], which reads
@@
We also established three analytic continuations of Saran’s function $F_{M}$. Notably, $F_{M}$ appears to have finer properties compared to Saran’s $F_{K}$ function. The fundamental properties of $F_{M}$ can be systematically derived by adapting techniques from prior studies on $F_{K}$ [6, 16, 17].
Finally, we would like to emphasize that all methods employed in this paper (see Table 1) are based on the series representations of multivariate special functions, rather than the integral representations. Among these methods, the uniformity approach currently has significant limitations (see [12, Section 7]). Now we propose three potential ways to overcome them:
@@ Acknowledgement
We wish to thank Malte Henkel for useful discussions with the first author.

### v2

@@ Abstract
We derive full asymptotic expansions for the Humbert function $\Phi_{1}$ in different limiting regimes of its variables. Our derivation employs various asymptotic methods and relies on key transformation formulae established by Erdélyi (1940), and Tuan and Kalla (1987). The efficiency of our asymptotic results is also illustrated through two applications: (1) analytic continuations of Saran’s function $F_{M}$, and (2) two limits arising in the study of the $1D$ Glauber-Ising model. Finally, some promising directions for future research are highlighted.
@@ 1.  Introduction
Applying the Euler transformation [19, Eq. (15.8.1)]
@@
Moreover, the generalized hypergeometric function ${}_{p}F_{q}$ is defined by [19, Equation (16.2.1)]
@@ 2.1.  Regime (i): $x\to\infty$
Using the connection formula for ${}_{2}F_{1}$ ([19, Eq. (15.8.2)])
@@ 2.3.  Regime (iii): $y\to\infty$
Recall the estimate [13, Eq. (3.4)]: as $t\to\pm\infty$,
@@
We close the path of integration by the semi-circle $\mathfrak{C}$, that is parameterized by $s=(N+1/2)\mathrm{e}^{\mathrm{i}\theta}$ $(|\theta|\leqslant\pi/2)$, and then we let $N\to\infty$ through integer values. Recall the estimate [13, p. 7]
@@ 2.4.  Regime (iv): $x\to\infty,\,y\to\infty$
and from asymptotics of the Humbert function $\Psi_{1}$ for two large arguments (see [12, Theorems 1.2 and 1.3]). We formulate the precise asymptotic expansions in the following two theorems.
@@
$\Phi_{1}={}$ $\frac{\Gamma(c)\Gamma(b-a)}{\Gamma(b)\Gamma(c-a)}A_{1}^{*}(x,y)+\frac{\Gamma(c)\Gamma(a-b)}{\Gamma(a)\Gamma(c-a)}A_{2}^{*}(x,y)+\frac{\Gamma(c)}{\Gamma(a)}A_{3}^{*}(x,y)$
$+\mathcal{O}\left(\left|y\right|^{-\Re(b)-w}\right)+\mathcal{O}\left(\left|y\right|^{\Re(a-b-c)-N}\mathrm{e}^{\Re(y)}\right),$
@@
$A_{2}^{*}(x,y)={}$ $\left(-\beta\right)^{b-a}\mathrm{e}^{-\beta}\left(1-x\right)^{-a}\sum_{k=0}^{M}\frac{\left(a-b\right)_{k}\left(a-c+1\right)_{k}}{k!}\,_{2}F_{2}\left[\begin{matrix}c-a,1-a-k\\ c-a-k,b-a+1-k\end{matrix};\beta\right]\left(-\beta\right)^{-k}\left(1-x\right)^{-k},$
$A_{3}^{*}(x,y)={}$ $\beta^{a-c}\left(1-x\right)^{a-b-c}\mathrm{e}^{y}\sum_{k=0}^{N-1}a_{k}^{*}\left(x,y\right)y^{-k},$
@@
2.5.  Regime (v): $x$ or $y$ small, $xy$ fixed
@@
It suffices to establish Assertion (1), as the proof of (2) follows analogously through application of the estimate in [12, Theorem 2.6(ii)]. Now assume that the conditions stated in (1) are valid.
@@
Estimate on $S_{2}$. For $n\geqslant N$, use the estimate [12, Theorem 2.5] to get
@@ 3.1.  Analytic continuations of $F_{M}$
Saran’s function $F_{M}$ is defined by ([20, Eq. (2.5)])
@@
which, in view of (1.3) and the estimate [13, Eq. (2.2)]
@@
which corrects Saran’s original form [21, p. 134, Eq. (4)]. Our results in Section 2.3 regarding asymptotics of $\Phi_{1}[y,zt]$ for large $t$ then provide a sufficient convergence condition for (3.3):
@@
The final continuation of $F_{M}$ follows from the following Mellin-Barnes integral, which is established by repeating the analysis in the proof of Theorem 2.3 and applying the asymptotic results of the Gauss hypergeometric function ${}_{2}F_{1}$ for large parameters in [13, Section 2.1].
@@ 3.2.  $1D$ Glauber-Ising model
which, by using the reduction formula [19, Eq. (15.4.3)]
@@ 4.  Discussion
which hold uniformly with respect to $n\in\mathbb{Z}_{\geqslant 0}$. More generally, it is necessary to extend the asymptotic expansions for ${}_{1}F_{1}[a;b;z]$ established in [25, Theorem 3.1], which reads
@@
We also established three analytic continuations of Saran’s function $F_{M}$. Notably, $F_{M}$ appears to have finer properties compared to Saran’s $F_{K}$ function. The fundamental properties of $F_{M}$ can be systematically derived by adapting techniques from prior studies on $F_{K}$ [16, 17]. Very recently, Dmytryshyn and his co-workers [6, 18] considered the problem of approximating $F_{M}$ functions and their ratios by branched continued fractions and also proposed some open problems. We anticipate that further properties of $F_{M}$ will be uncovered.
Finally, we would like to emphasize that all methods employed in this paper (see Table 1) are based on the series representations of multivariate special functions, rather than the integral representations. Among these methods, the uniformity approach currently has significant limitations (see [11, Section 7]). Now we propose three potential ways to overcome them:
@@ Acknowledgement
We would like to thank Malte Henkel and Roman Dmytryshyn for many helpful discussions with the first author. The first author was supported by the Fundamental Research Funds for the Central Universities (CUSF-DH-T-2025025).